OpenStudy (anonymous):

\[a_{n+2}=\frac{a_n}{(n+1)(n+2)}\] \[a_9=\frac{a_7}{8\cdot9}=\frac{a_0}{2\cdot3\cdot4\cdot5\cdot6\cdot7\cdot8\cdot9}\]

OpenStudy (anonymous):

\[y=\sum_{\infty}^{n=0}\]

OpenStudy (anonymous):

....

OpenStudy (anonymous):

@phi please help

OpenStudy (phi):

after some thinking I decided that \( a_n= \frac{a_0}{n!} \) and you want \[\sum_{n=0}^{\infty}\frac{a_0}{n!}= a_0\sum_{n=0}^{\infty}\frac{1}{n!}\] See http://www.khanacademy.org/math/calculus/integral-calculus/v/approximating-functions-with-polynomials--part-3